Please use hand work and formulas, as I cannot use excel and it helps me better understand.

Problem 1 What are the present value and future value of $50 to be received at the BEGINNING of each year for the next 5 years if the discount/compounding is 11%?
Answer #1

Present value of annuity received at the beginning of each year = Amount + Amount*PVAF(r%, n-1 periods)

Hence, present value of $50 to be received at the BEGINNING of each year for the next 5 years

= 50 + 50*PVAF(11%, 4 years)

= 50 + 50*3.102

= $205.1

Future Value of annuity received at the beginning of each year = (1+r)*P[{(1+r)n-1}/r]

= (1+0.11)*50[{(1+0.11)5-1}/0.11]

= $345.64